Murphy's Rebar Company makes safety vests for rebar workers, and this is the only product they manufacture. They have provided the information below concerning their most recent month of operation:

Number of units produced 14,600
Variable costs per unit:
Direct materials $ 137
Direct labor $ 75
Variable manufacturing overhead $ 4
Variable selling and administrative expenses $ 11
Fixed costs:
Fixed manufacturing overhead $ 846,800
Fixed selling and administrative expenses $ 233,600

Murphy's had NO beginning or ending inventories.

Required:

a. Compute the unit product cost for their vests under absorption costing.

b. Compute the unit product cost for their vests under variable costing.
Answer #1

a.Fixed manufacturing overhead per unit=846,800/14,600

=$58

Hence unit product cost=Direct materials+Direct labor+Variable manufacturing overhead+Fixed manufacturing overhead

=137+75+4+58

=$274 per unit

b.unit product cost=Direct materials+Direct labor+Variable manufacturing overhead

=137+75+4

=$216 per unit